A solid in the shape of a hemisphere with a radius of 2 units, has its base in the xy-plane
and the centre of the base at the origin. If the density of the solid is given by the function
ρ(x, y,z) = xyz, determine the mass of the hemisphere.
1
Expert's answer
2019-03-08T09:00:04-0500

Mass of the hemisphere could be calculated as:

"M=\int\rho\left(x,y,z\right)dV"

In polar coordinates:

"x=rcos\varphi sin\theta,\ y=rsin\varphi sin\theta,\ \ z=rcos\theta"

This gives us:


"M=\int\rho\left(x,y,z\right)dV=\int_{0}^{2}\int_{0}^{2\pi}\int_{0}^{\pi/2}\left(rcos\varphi sin\theta\right)\left(rsin\varphi sin\theta\right)\left(rcos\theta\right)r^2sin\theta d\theta d\varphi dr="

"=\int_{0}^{2}{r^5dr\int_{0}^{2\pi}{cos\varphi sin\varphi d\varphi\int_{0}^{\pi/2}{sin^3\theta cos\theta}}}d\theta=(\frac{64}{6})(0)(0.25)=0"

So, the total mass of the hemisphere is zero